a wire 10 m long is perpendicular to magnetic field of induction 1 T.calculate the force of the wire when it carries a current of 1 A

Grade:10

1 Answers

Pragya Verma
3 years ago
we know force on wire is given by F=BIL sin \cap, here angle is 90 , so sin\large \sin \Theta =1
,length of wire is 10 m , magnetic feild B -= 1T , threfore putting values 
F =1T*10m*1A*1=10 N , as 1T =N/A.m
so force on wire is 10 N